### v1

# An Early Look at the Performance of IGRINS–2 at Gemini-North with Application to the ultrahot Jupiter, WASP-33 b

Abstract
Ground-based high-resolution spectroscopy enables precise molecular detections, and velocity-resolved atmospheric dynamics, offering a distinct advantage over low-resolution methods for exoplanetary atmospheric studies. IGRINS-2, the successor to IGRINS, features improved throughput and enhanced sensitivity to carbon monoxide by shifting its K-band coverage by 36 nm to longer wavelengths. To evaluate its performance, we attempt to investigate the atmospheric characteristics of WASP-33 b. Observations were conducted on 2024 January 7 for a total of 2.43 hours; This includes 1.46 hours in the pre-eclipse phase to capture the planet’s thermal emission spectrum. Our analysis, employing a composite atmospheric model, confirms the presence of a thermally inverted atmosphere on the planet’s dayside with a signal-to-noise ratio (SNR) of 7.5. More specifically, we capture CO, H2O, and OH with SNRs of 5.2, 4.2, and 4.4, respectively. These results are consistent with previous studies and demonstrate that IGRINS-2 is well-suited for detailed investigation of exoplanetary atmospheres. We anticipate that future observations with IGRINS-2 will further advance our understanding of exoplanetary atmospheres.
1 Introduction
High-resolution infrared spectroscopy (R 15,000) has become a cornerstone for characterizing exoplanetary atmospheres, offering unique advantages over low-resolution spectroscopy. Unlike low-resolution spectroscopy, which primarily probes deeper layers of an atmosphere, high-resolution spectroscopy enables the detection of molecular species across a broader range of pressures, providing insights into thermal profiles and atmospheric dynamics (Snellen et al., 2010). Additionally, high-resolution spectroscopy is highly sensitive to velocity-resolved information, allowing for the study of wind patterns and rotational dynamics (Wardenier et al., 2024). These capabilities offer a deeper understanding of atmospheric chemistry and planetary formation histories (Öberg et al., 2011; Booth et al., 2017; Chachan et al., 2023). High-resolution cross-correlation spectroscopy (HRCCS), in particular, has become an essential tool for exploring exoplanetary atmospheres. By aligning observed spectra with theoretical templates, HRCCS allows for the detection of molecular and atomic species and provides insights into thermal and velocity structures. This technique has been successfully applied to various orbital geometries (i.e. transmission and thermal emission), enabling a comprehensive investigation of exoplanetary atmospheres (Snellen et al., 2010; de Kok et al., 2013; Brogi et al., 2023).
The Immersion GRating INfrared Spectrograph 2 (IGRINS-2; Lee et al., 2022; Oh et al., 2024) has been recently integrated as a facility instrument at the 8.1 m Gemini-North Observatory. IGRINS-2 builds upon the success of its predecessor IGRINS (Yuk et al., 2010; Park et al., 2014; Mace et al., 2016, 2018) by offering high-resolution spectroscopy in the near-infrared H (1.442–1.836 µm, R 50,000) and K (1.892–2.554 µm, R 45,000) bands. With these specifications, the instrument has a particular advantage for HRCCS in probing the chemical composition of hot Jupiters due to its wide simultaneous wavelength coverage, as demonstrated by previous high-resolution studies using IGRINS (e.g., Line et al., 2021), Gemini/MAROON-X (e.g., Pelletier et al., 2023), VLT/ESPRESSO (e.g., Borsa et al., 2021), VLT/CRIRES+ (e.g., Yan et al., 2023), and TNG/GIANO (e.g., Giacobbe et al., 2021). Additionally, with IGRINS-2 now operating in the northern hemisphere, we not only gain access to a new population of northern targets, which have not been explored by IGRINS at Gemini-South111IGRINS was operated at Gemini-South as a visiting instrument until April 2024 and is now in operation at McDonald Observatory, but also open possibilities for complementary follow-up observations using other high-resolution spectrographs such as Gemini-North/MAROON-X (Seifahrt et al., 2018), LBT/PEPSI (Strassmeier et al., 2015), Subaru/IRD (Tamura et al., 2012), Subaru/HDS (Noguchi et al., 2002), and Keck I/HIRES (Vogt et al., 1994), Keck II/NIRSPEC (McLean et al., 1998; Martin et al., 2018).
WASP-33 b, an ultrahot Jupiter with an equilibrium temperature of approximately 2781 K (Chakrabarty & Sengupta, 2019), serves as an ideal target for testing IGRINS-2’s capabilities in HRCCS due to its brightness and high equilibrium temperature. Planets in this temperature regime have been investigated with high-resolution spectroscopy in both the infrared and optical wavelengths, revealing a variety of atomic and molecular species such as CO, H2O, OH, and Fe (e.g., Yan et al., 2020; Cabot et al., 2021; Brogi et al., 2023; Ramkumar et al., 2023; Damasceno et al., 2024). The WASP-33 system is a well-studied system with key properties including a short orbital period of 1.22 days (Chakrabarty & Sengupta, 2019) and orbiting a bright A5 star (, ). Discovered by Christian et al. (2006) through SuperWASP transit photometry, WASP-33 b orbits its host star in a near-polar retrograde orbit, likely influenced by dynamical interactions such as the Kozai-Lidov effect (Naoz et al., 2011).
Previous high-resolution spectroscopy studies have revealed a thermally inverted atmosphere on the dayside of WASP-33 b, characterized by significant photodissociation of H2O and the detection of various molecules and atomic species. Emission spectroscopy has identified CO, OH, and Fe I as prominent components in the dayside atmosphere, while transmission spectroscopy has detected AlO and Si I in the terminator region (Nugroho et al., 2021; Cont et al., 2021, 2022a, 2022b; Yan et al., 2022; van Sluijs et al., 2023a). These findings suggest a complex atmospheric structure, with early inferences pointing to a thermal inversion on the dayside. Yang et al. (2024) extend this understanding by suggesting that thermal inversion also occurs in the terminator region, offering new insights into the atmospheric dynamics and chemical processes of WASP-33 b.
The recent study by Finnerty et al. (2023, hereafter, F23) using the Keck Planet Imager and Characterizer (KPIC, R 35,000) provided comprehensive insights into the atmospheric composition of WASP-33 b. Their analysis validated a thermally inverted dayside atmosphere with significant detections of CO, H2O, and OH in emissions. Following the detection of molecular species, Bayesian atmospheric retrievals (Brogi & Line, 2019) were performed to place quantitative constraints on the composition of WASP-33 b’s dayside atmosphere. The model fits indicated a carbon- and metal-enriched environment, with a gas-phase C/O ratio of 0.80.2 and evidence for a metal-enriched atmosphere (2–15 solar). These findings are consistent with formation scenarios that involve accretion near the CO or CO2 snow lines (Öberg et al., 2011) or enrichment by carbon-rich grains interior to the H2O snowline (Chachan et al., 2023), as suggested by F23. Additionally, F23 found tentative evidence for a 12CO/13CO ratio of 50 consistent with values expected in protoplanetary disks.
In this paper, we present an early look at the performance of IGRINS-2 for the atmospheric characterization of an exoplanet using WASP-33 b. Our findings demonstrate that IGRINS-2 is capable of detecting molecular signatures, showcasing its potential for future exoplanetary atmospheric studies. We provide a detailed account of our observational methods, data reduction techniques, and analysis, illustrating the instrument’s suitability for high-resolution spectroscopic investigations.
2 Upgrades on IGRINS-2
The upgrades from IGRINS to IGRINS-2 include enhanced sensitivity and a shift in spectral coverage, which are critical for HRCCS of exoplanetary atmospheres. This shift resulted from the process of aligning the echellogram on the detector by adjusting the main disperser, the immersion grating. The shift toward longer wavelengths was not intentional; rather, the wavelength range was determined while optimizing optical performance, avoiding bad pixels, etc., within the required K-band wavelength range of 1.96–-2.46 µm. To isolate instrumental differences from site-specific effects (e.g., telescope throughput, atmospheric transmission), we compared flat lamp frames from the calibration systems of both instruments. The major changes in optics include replacing the mirror of the IGRINS’s calibration system with a beam splitter in IGRINS-2 and using a dichroic mirror with enhanced response at the edges of each channel in the spectrograph. We corrected these differences by interpolating and normalizing the response function based on the transmittance of each optical component, though these corrections had an insignificant impact on the final spectral response (see Figure 1).
2.1 K-Band Spectral Shift
One of the most significant changes in IGRINS-2 is the shifts of the H and K-band spectra to longer wavelengths. Specifically, the K band now extends approximately 36 nm further to longer wavelengths compared to IGRINS. This shift could be crucial because it captures additional spectral features, particularly CO molecular lines that begin to appear around 2.3 µm. By covering more CO lines in the K band, as shown in the bottom right panel of Figure 1, IGRINS-2 improves the detection and characterization of molecular features in exoplanetary atmospheres, resulting in an increased signal in the cross-correlation analysis.
2.2 Sensitivity Enhancements
Another significant advancement in IGRINS-2 is the improvement in the quantum efficiency (QE) of the detector in the K band. The QE of IGRINS in the K band was approximately 80%–85% (Yuk et al., 2010), whereas IGRINS-2 achieved an average QE of 96% (C. Park et al. in preparation). This increase in QE results in more efficient photon collection, particularly at longer wavelengths, enhancing the overall sensitivity of the instrument and improving the detection of weaker molecular features.
The impact of this enhanced QE is evident in the noticeable increase in response functions obtained in the extended portion of the K band (see the bottom right panel of Figure. 1). This improvement in sensitivity is crucial for detecting faint molecular signals in exoplanet atmospheres, as the planet’s signal-to-noise ratio is directly proportional to the stellar signal-to-noise ratio (SNR), as described by Birkby (2018). Together, these advancements, particularly in the longer wavelength region, should enable more robust atmospheric detections, especially for species such as CO, which are prominent in the K band.
3 Observations and Data reduction
3.1 Observations
On UT date 2024 January 7, we observed WASP-33 b as part of the test observations of the IGRINS-2 attached to the Gemini-North telescope. The observations of WASP-33 b were designed to probe the dayside thermal emission spectrum and benchmark its atmospheric characterization capabilities, evaluating the effectiveness of IGRINS-2 for high-resolution exoplanet spectroscopy. We obtained a continuous sequence of 48 AB pairs over 2.43 hours covering the pre-eclipse orbital phases (0.399 0.481) with the secondary eclipse of WASP-33 b beginning at orbital phase 0.451. The integrated time for a single AB pair was 80 seconds.
Figure 2 shows the SNR, seeing, and airmass variations during the observations. Throughout the observation period, the relative humidity remained stable at approximately 9%, as recorded by the Maunakea Weather Center111http://mkwc.ifa.hawaii.edu/archive/ at the Canada–France–Hawaii Telescope station, which is located near Gemini-North. Seeing, obtained from the Differential Image Motion Monitor (DIMM), varied between 0″.52 and 0″.97, with an average seeing of approximately 0″.68.
3.2 Data reduction
The data reduction was performed using the IGRINS Pipeline Package Version 3.0 (PLP; Kaplan et al., 2024), which was initially created for the original IGRINS. The igrins2-dev branch of the pipeline, which contains the necessary modifications for IGRINS-2, was utilized. Flat-fielding was applied to the raw data to adjust for pixel-to-pixel differences. Bad pixels were identified and removed from both flat-off and flat-on images. One-dimensional spectra were extracted using a modified version of Horne (1986) optimal extraction techniques. The final result consists of one-dimensional spectra for each diffraction order.
After initial data reduction, we used the IGRINStransit (Mansfield & Lee, 2024) package for secondary wavelength calibration and barycentric-velocity correction which was described in Line et al. (2021). IGRINStransit also includes routines for the data cube production, pixel trimming, and exclusion of bad orders. We trimmed 100 pixels from the edges of each order due to their low SNRs and discarded bad orders at the edge of each band that were strongly contaminated by telluric absorption lines during calibration. As a result, five orders were discarded out of the total 53 orders: the longest order (2.520–2.554 µm) and the shortest three orders (1.866–1.932 µm) in the K band, as well as the shortest order (1.441–1.461 µm) in the H band. Each remaining order contains 1848 pixels.
To minimize noise from stellar variability and systematics, we cropped the data following the secondary eclipse phase before analysis. The complete dataset (including eclipse observations) was preserved for trail plot visualization to maintain full orbital phase context.
The data cube, produced by stacking the time series spectra for each order and pixel, contains a mix of the stellar spectrum, telluric absorption, planet signal, and photon noise. To extract the planet signal, it is crucial to isolate and remove the stationary stellar and telluric lines, which remain quasi-stationary during observations, while the planet’s spectrum shifts due to its orbital motion. This can be achieved through singular value decomposition (SVD), which decomposes the time-series data into principal components (see Birkby et al., 2013; de Kok et al., 2013; Giacobbe et al., 2021; Line et al., 2021). The dominant stationary features are removed, leaving the planet signal in the residuals. The optimal number of SVD components to remove can vary depending on observing conditions and the level of telluric absorption. Following the approach of F23, we found that using four singular values (SVs) yielded the highest SNR via a cross-correlation, though using three or five SVs resulted in minimal changes within one .
3.3 Wavelength calibration stability
Wavelength calibration stability is critical to the success of HRCCS, as even minor instabilities in the wavelength solution can adversely affect the alignment between observed spectra and atmospheric models during the cross-correlation process. We evaluated the wavelength calibration stability using two steps: First, due to the broad stellar absorption lines caused by the rapid rotation of the host star, the telluric lines were primarily used to evaluate the stability of the initial wavelength solution. Then, a secondary wavelength calibration using scipy.curve_fit was applied to refine the wavelength solution. This corrected any residual shifts in the wavelength calibration across the spectral orders. Figure 3 shows the resulting wavelength calibration shifts over time for both the H & K bands, where each point represents the shift for an individual spectral order.
The analysis demonstrates that these two calibration steps achieved high stability, with wavelength shifts maintained within the sub-pixel range throughout the observations. This level of stability ensures that the Doppler shifts due to planetary motion can be accurately captured during the HRCCS process, minimizing errors in molecular signal detection.
4 Molecular Signal Detection via Cross-correlation
4.1 Atmospheric models
To perform cross-correlation analysis, we generated thermal emission spectra for WASP-33 b with a resolving power of 250,000 using the petitRADTRANS package (Mollière et al., 2019, 2020). The atmospheric structure is modeled across a pressure range from 10-6 to 102 bar, divided into 100 layers with uniform spacing in logarithmic pressure. Models for each molecule and pressure-temperature (P-T) profile used in the analysis are shown in Figure 4. Consistent with several previous studies, we assumed a thermal inversion profile for the dayside atmosphere (Nugroho et al., 2021; van Sluijs et al., 2023b; Yan et al., 2022; Finnerty et al., 2023) and we adopted Guillot P-T profile (Guillot, 2010). Molecular species included in the models were H2O (Rothman et al., 2010), CO (Li et al., 2015), and OH (Gordon et al., 2022) from HITRAN/HITEMP, chosen for following the results from F23. These models were built based on key parameters from F23, including reference pressure, mass mixing ratios (MMR), and the Guillot P-T profile parameters , and IR (Finnerty et al., 2023, private communication). Stellar and planetary properties were adopted from Chakrabarty & Sengupta (2019), and a PHOENIX model was used for the stellar template (Husser et al., 2013). The parameters used for the atmospheric models are summarized in Table 1.
| Parameter | Value |
| Stellar effective temperature | 7430 K a |
| Stellar mass | 1.495 a |
| Stellar radius | 1.444 a |
| Planetary mass | 2.093 a |
| Planetary radius | 1.593 a |
| Orbital period | 1.21987 days a |
| Transit duration | 2.854 hours a |
| Kp | b |
| Top-of-atmosphere temperature | 3200 K |
| Intrinsic temperature | 50 K |
| log MMRCO | c |
| log MMR | c |
| log MMROH | c |
| log | 0.5 c |
| log IR | c |
| CO ratio | 0.8 c |
| log reference pressure | bar |
4.2 Cross-correlation maps
To validate the performance of IGRINS-2 on HRCCS, we cross-correlated the post-SVD residual spectra with atmospheric models. These models were convolved with the IGRINS-2 instrumental profile to match the resolving power of the observations (). The cross-correlation function (CCF) was computed by systematically shifting the model spectra across a range of radial velocities and calculating the correlation at each step. We followed the same approach described in Brogi & Line (2019) and Line et al. (2021) that is used to detect the atmospheric signal using IGRINS (see Brogi et al., 2023; Kanumalla et al., 2024; Smith et al., 2024; Weiner Mansfield et al., 2024). Frames taken during the secondary eclipse, when the planet is hidden behind its host star, were not included when generating - maps.
Figure 5 presents the resulting cross-correlation maps, generated using the composite atmospheric model with CO, H2O, and OH. A planetary trail before ingress is clearly shown in the top panel. The dashed red line corresponds to the Doppler shift induced by the planet’s orbital motion. The alignment of this trail with the predicted velocity curve, derived from WASP-33 b’s orbital parameters, confirms the detection of the planet’s atmospheric signal. The cross-correlation map (bottom panel in Fig. 5) shows the composite atmospheric thermal emission with SNR = 7.5. The dashed white lines indicate the expected planet’s radial velocity () from F23 and a system velocity () from Gontcharov (2006). SNR is calculated following the same approach outlined in Weiner Mansfield et al. (2024) using the 3 clipping method. The positive cross-correlation value further suggests that the detected spectral lines are in emission rather than absorption, supporting the presence of a thermally inverted dayside atmosphere.
We also attempted to detect individual molecular signals from CO, H2O, and OH in the atmosphere using models in Figure 4. These results are shown in Figure 6, formatted similarly to the result of the composite atmospheric model. CO is robustly detected at the expected planetary velocity, consistent with previous studies and confirming its presence in the planet’s dayside emission. The cross-correlation results of H2O and OH exhibit velocity shifts, which may suggest additional atmospheric processes at play, as individual molecular species have shown similar shifts in previous studies (Brogi et al., 2023; Cont et al., 2021, 2024). One possibility is that the intense dayside heating leads to the thermal dissociation of H2O into H and OH, resulting in non-uniform molecular distributions across the atmosphere (e.g., Lothringer et al., 2018; Parmentier et al., 2018; Nugroho et al., 2021, Finnerty et al., in review). Alternatively, strong atmospheric dynamics, such as high-altitude winds or circulation patterns, may contribute to the observed Doppler shifts by redistributing molecular species across different regions of the atmosphere. Follow-up observations, particularly those covering a broader orbital phase range, will be essential to confirm these velocity shifts and further investigate the planet’s atmospheric circulation and chemistry.
Notably, we achieved high SNR molecular detections on the planetary signal (0.05 orbital phase) with an observation time of only 1.46 hours. This is shorter than the typical observation time of 2–4 hours required in similar studies conducted with 8-m class telescopes (Nugroho et al., 2021; Line et al., 2021; Brogi et al., 2023; Finnerty et al., 2023). This result highlights the sensitivity and efficiency of IGRINS-2 for HRCCS.
5 Summary
In this paper, we presented an early look at the performance of IGRINS-2 with HRCCS to study the dayside atmosphere of the ultrahot Jupiter WASP-33 b. This assessment focused on the instrument’s detectability of molecular species, such as CO, H2O, and OH.
We successfully detected clear cross-correlation signals from these molecular species in the dayside atmosphere of WASP-33 b with a combined SNR of 7.5. The detection was achieved by cross-correlating the post-SVD residual spectra with atmospheric models convolved to match the instrumental resolution of R=45,000. The results confirmed the presence of molecular species that were consistent with previous results from other high-resolution spectroscopy studies (e.g., Nugroho et al., 2021; Cont et al., 2021, 2022a, 2022b; Yan et al., 2022; Finnerty et al., 2023), supporting the conclusion of a thermally inverted atmosphere for WASP-33 b. The detected planetary trail, which aligns with the expected radial velocity curve, further corroborates the planetary origin of the signal and underscores the precision of IGRINS-2 in tracking Doppler-shifted atmospheric features.
The cross-correlation maps demonstrate the strength of the planetary signal and its alignment with the predicted radial velocity of the planet. This not only validates IGRINS-2’s performance for HRCCS but also highlights its potential for atmospheric characterization of exoplanets. The clear visibility of the planetary trail during the pre-eclipse orbital phases demonstrates the capability of IGRINS-2 to detect molecular features in exoplanetary atmospheres with high sensitivity.
Our results are in good agreement with previous studies on ultrahot Jupiters (e.g., Finnerty et al., 2023; Brogi et al., 2023), further supporting the reliability of atmospheric models for these extreme environments. In particular, our detection of CO and H2O is consistent with theoretical predictions for high-temperature atmospheres where strong molecular excitation occurs (Madhusudhan et al., 2011). The presence of OH in the atmosphere adds to the growing body of evidence for complex chemical processes at play in these planetary systems, particularly in atmospheres experiencing significant photodissociation and thermal inversion.
In conclusion, IGRINS-2 has demonstrated its ability to perform precise HRCCS on exoplanets, providing valuable insights into the atmospheric composition of WASP-33 b. This study highlights the instrument’s effectiveness in detecting molecular signatures in the thermal emission spectra of ultrahot Jupiters and establishes IGRINS-2 as a promising tool for future exoplanet studies. Further observations across different planetary systems will be necessary to explore the broader applicability of IGRINS-2 in atmospheric characterization and to extend our understanding of the diverse chemical environments in exoplanet atmospheres.
Acknowledgments
Y.H.C. thanks Stefan Pelletier, Woowon Byun, and Jaerim Koo for improving the manuscript. This work used the Immersion Grating Infrared Spectrograph 2 (IGRINS-2) developed and built by a collaboration between Korea Astronomy and Space Science Institute (KASI) and the International Gemini Observatory. Based on observations obtained at the international Gemini Observatory, a program of NSF NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation on behalf of the Gemini Observatory partnership: the U.S. National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). This work was enabled by observations made from the Gemini North telescope, located within the Maunakea Science Reserve and adjacent to the summit of Maunakea. We are grateful for the privilege of observing the Universe from a place that is unique in both its astronomical quality and its cultural significance.